Flamo:

Taking the square root of which of the following expressions gives the length of the line segment shown?

Flamo:

1 attachment
Flamo:

Same thing?

Angle:

mhm

Flamo:

*sigh*....

Flamo:

\[d = \sqrt{(-5-2)^2 + (3- (-1)})^2 \]

Flamo:

?

Angle:

perfect ^_^

Flamo:

\[d = \sqrt{(-7)^2 - (4)^2 }\]

Flamo:

That's the answer.

Angle:

how did you get a minus before the 4?

Flamo:

Since there is a "3 -" Then the Plus Sign leaves, and replaces it with a Minus

Flamo:

I think.

Angle:

3 - (-1) = 3 + 1 = 4 no minus signs?

Flamo:

o

Flamo:

XD Idk how I got that. Anyways I have to go now.

Flamo:

Thank you for helping me!

Angle:

aye, good luck